The following information is provided to assist you in evaluating the performance of the production operations of Studio Company:

Units produced (actual) 46,000
Master production budget
Direct materials $131,010
Direct labor 111,160
Overhead 166,740
Standard costs per unit
Direct materials $1.65 × 2 gallons per unit of output
Direct labor $14 per hour × 0.2 hour per unit
Variable overhead $12.50 per direct labor-hour
Actual costs
Direct materials purchased and used $140,785 (76,100 gallons)
Direct labor 118,548 (8,880 hours)
Overhead 166,200 (61% is variable)

Variable overhead is applied on the basis of direct labor-hours.

Required:

Calculate all variable production cost price and efficiency variances and fixed production cost price and production volume variances. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)
Answer #1

Price variance

Efficiency variance

Direct materials

$15220 U

$26235 F

Direct labor

$5772 F

$4480 F

Variable overhead

$9618 F

$4000 F

Fixed overhead

$2672 F

$10710 F

Direct materials price variance = AQ*(AP-SP) = 76100*((140785/76100)-1.65) = $15220 U

Direct materials efficiency variance = 1.65*(76100-(46000*2)) = -26235 = $26235 F

Direct labor price variance = AH*(AR-SR) = 8880*((118548/8880)-14) = -5772 = $5772 F

Direct labor efficiency variance = 14*(8880-(46000*0.2)) = -4480 = $4480 F

Variable overhead price variance = AH*(AR-SR) = 8880*(((166200*61%)/8880)-12.50) = -9618 = $9618 F

Variable overhead efficiency variance = 12.50*(8880-(46000*0.2)) = -4000 = $4000 F

Fixed overhead price variance = Actual fixed overhead - budgeted fixed overhead = 64818-67490 = -2672 = $2672 F

Budgeted units = budgeted materials cost / standard direct materials cost per unit = 131010/(1.65*2) = $39700

Budgeted fixed overhead = 166740-(12.50*0.2*39700) = $67490

Actual fixed overhead = 166200*39% = $64818

Fixed efficiency (Production volume) variance = budgeted fixed overhead – applied fixed overhead = 67490 – ((67490/39700)*46000) = -10710 = $10710 F